# Data Selection via Optimal Control for Language Models

### v1

@@ Abstract
This work investigates the selection of high-quality pre-training data from massive corpora to enhance LMs’ capabilities for downstream usage. We formulate data selection as a generalized Optimal Control problem, which can be solved theoretically by Pontryagin’s Maximum Principle (PMP), yielding a set of necessary conditions that characterize the relationship between optimal data selection and LM training dynamics. Based on these theoretical results, we introduce PMP-based Data Selection (PDS), a framework that approximates optimal data selection by solving the PMP conditions. In our experiments, we adopt PDS to select data from CommmonCrawl and show that the PDS-selected corpus accelerates the learning of LMs and constantly boosts their performance on a wide range of downstream tasks across various model sizes. Moreover, the benefits of PDS extend to ~400B models trained on ~10T tokens, as evidenced by the extrapolation of the test loss curves according to the Scaling Laws. PDS also improves data utilization when the pre-training data is limited, by reducing the data demand by 1.8 times, which mitigates the quick exhaustion of available web-crawled corpora. Our code, data, and model checkpoints can be found in https://github.com/microsoft/LMOps/tree/main/data_selection.
@@
Figure 1: Scaling curves of average accuracy on 9 widely-used downstream tasks with respect to computation (a) and model sizes (b). We select pre-training corpora from the CommonCrawl and pre-train LMs on the selected data. PDS is compared with the Redpajama data cleaning pipeline [71]. The computation curves are calculated based on the training of a 1.7B LM.
@@ 1 Introduction
With the thriving of language models (LMs; 38, 10), the role of data selection for pre-training becomes increasingly important, which aims at identifying valuable pre-training instances to accelerate model learning or improve downstream performance [4]. This focus enables researchers to explore the limit of LMs in the face of increasing training data demands [12, 60, 67]. It also helps to reduce the computational costs during pre-training [66], and addresses the potential limitations caused by available Internet data [74, 59]. Without doubt, pre-training data selection is highly valuable for both research and industry sectors.
@@
Unlike previous works relying primarily on manually crafted heuristics [70, 79], we connect data selection with classical Optimal Control theory [47], where control variables in a dynamic system are optimized to achieve desired objectives. This mathematical formulation allows fine-grained white-box analysis of how the control variables drive a dynamic system from one state to another. In particular, by treating data selection as the control variables (i.e., whether a data point is included in pre-training), the LM pre-training process as the dynamic system, and the LM’s downstream performance as the objective, we leverage Pontryagin’s Maximum Principle (PMP; 63) to derive the necessary conditions for optimal data selection in theory. These results offer a rigorous, theory-driven alternative to the ad-hoc trial-and-error practices that currently dominate LM pre-training.
@@
Based on our theoretical results, we introduce PMP-based Data Selection (PDS), a framework that selects high-quality pre-training data at scale, by solving the equation system induced by the PMP conditions. Balancing effectiveness and efficiency, PDS first solves the equation system for the optimal data selection on a proxy dataset (e.g., 0.2B tokens), assigning a quality score to each instance based on its impact on downstream tasks. After that, a data scorer (e.g., 125M) is trained to predict the quality scores and then infers scores on the target corpus (e.g., 50B tokens). Finally, the predicted scores guide data selection for pre-training LMs with various sizes (e.g., 1.7B) from scratch.
@@
Unlike previous pre-training data selection methods based on deduplication [70, 1], pattern matching [79], or single checkpoint performance [22, 43], which are agnostic to the pre-training process of LMs, PDS exploits the highly dynamic nature of LM pre-training through the theoretical optimal control perspective. On the other hand, compared to methods that incorporate signals from the LM training process in an online fashion [81, 75], PDS operates offline, before the training begins, which avoids additional training-time computation overhead and allows for training LMs with arbitrary configurations while performing PDS only once. Furthermore, PDS only filters the training corpus, leaving highly optimized pre-training pipelines largely intact. Most importantly, PDS enjoys a strong theoretical basis, opening up the black box of understanding the impact of individual data points on LM pre-training.
@@
In our experiments, we select data from the CommonCrawl [^1] https://commoncrawl.org/ with PDS and pre-train LMs with 160M, 470M, 1B, and 1.7B parameters from scratch. We observe around 2 times speed-up in pre-training on the 1.7B LM and constant improvement in downstream tasks and language modeling performance across all model sizes compared to state-of-the-art baselines. Extrapolating these results using the Scaling Law [44, 40], we show that the benefits remain consistent for ~400B LMs trained on ~15T tokens. Besides, PDS enhances data utilization in a data-constrained setting, reducing the pre-training data demand by 1.8 times, which is a critical advantage as the LM community is running out of data [74]. We also conduct extensive analysis and ablation studies on the key factors of PDS to facilitate further research on data selection.
@@ 2.1 Problem Formulation
We study an LM parameterized with ${\bm{\theta}}\in{\mathbb{R}}^{N}$, pre-trained from scratch on a dataset $\mathcal{D}=\{x_{n}\}_{n=1}^{|\mathcal{D}|}$, over a total $T$ steps. Data selection [79, 22] aims at identifying a subset $\mathcal{D}^{\prime}$ from $\mathcal{D}$, such that LMs trained on $\mathcal{D}^{\prime}$ achieve better downstream performance, measured by a lower downstream loss $J({\bm{\theta}})$.
The pre-training process renders $J({\bm{\theta}})$ as a function of $\mathcal{D}^{\prime}$, which can be fully characterized by a data quality score vector ${\bm{\gamma}}=\left[\gamma_{1},\gamma_{2},\cdots,\gamma_{|\mathcal{D}|}\right]^{\top}$ in a $|\mathcal{D}|$-dimensional simplex $U$, where $U=\left\{\left.[\gamma_{1},\gamma_{2},\cdots,\gamma_{|\mathcal{D}|}]^{\top}\right|\sum_{n=1}^{|\mathcal{D}|}\gamma_{n}=1\text{ and }\gamma_{n}\geq 0\text{ for }1\leq n\leq{|\mathcal{D}|}\right\}$  [^2] Only the relative data quality is meaningful for data selection. Therefore, we ensure the sum of the quality scores to 1 to avoid the impact of their individual scales.. A higher quality score in ${\bm{\gamma}}$ indicates the corresponding instance is more helpful to reduce $J({\bm{\theta}})$, and thus the LM should learn more from the instance. This results in the following general pre-training loss, defined as the weighted sum of the per-instance loss by ${\bm{\gamma}}$:
@@
where $l(x_{n},{\bm{\theta}})=-\log p_{{\bm{\theta}}}(x_{n})$. The goal of data selection is thus to find ${\bm{\gamma}}$ that reduces the downstream loss $J({\bm{\theta}})$, and then select instances with the highest scores in ${\bm{\gamma}}$. For simplicity, we assume that the LM is trained using Gradient Decent (GD) for $0\leq t<T$, with the derivation under the Adam optimizer [45] provided in Appendix C:
@@
Motivated by the literature on learned optimizers [56], we optimize ${\bm{\gamma}}$ by minimizing the area under the curve (AUC; 18) of $J({\bm{\theta}}_{t})$, which is approximated by the cumulative sum of $J({\bm{\theta}}_{t})$ over the pre-training process:
@@
Intuitively, a lower AUC corresponds to faster convergence of the loss and improved final downstream performance. Unlike evaluating $J({\bm{\theta}}_{t})$ at individual time steps, the AUC captures the overall LM training dynamics. As shown in Appendix A, minimizing the AUC essentially enhances the constants in the LM’s Scaling Laws [44], leading to substantial improvements in LM learning.
@@ 2.2 Data Selection as Optimal Control
We recognize the optimization problem in Eq. (3) is analogous to a discrete time optimal control problem [47], where $J(\cdot)$ is the cost function, the model parameters ${\bm{\theta}}$ are the state variables evolving according to Eq. (2), and the data quality scores ${\bm{\gamma}}$ are the control variables to be optimized within $U$. This perspective provides a convenient framework to formulate the pre-training data selection problem from a more theoretical perspective.
@@
Optimal control problems can be solved by a powerful tool known as Pontryagin’s Maximum Principle (PMP; 63), which provides a set of necessary conditions for the optimal control variables and their corresponding state variables (See Appendix B for its formal expression). However, standard PMP conditions allow the optimal control to vary over time, whereas in Eq. (3), the control variables ${\bm{\gamma}}$ are constrained to be time-invariant due to the offline nature of data selection in our setting. This typically makes the optimization problem more challenging due to the shrinking of feasible region. In the following, we present the PMP conditions for data selection under this constraint:
@@
We prove Theorem 2.1 in Appendix B using the standard PMP conditions and the Lagrange multiplier method, and provide an illustration for this theorem in Figure 2. By inspecting the PMP conditions for data selection, we can see that Eq. (4) ensures the LM parameters, trained with the optimal data quality scores, continue to evolve via GD. As illustrated in Figure 2 (Left), Eq. (5) defines ${\bm{\lambda}}^{*}_{t}$, a “target vector” suggesting the ideal gradient direction formed only by high-quality data points. In particular, ${\bm{\lambda}}^{*}_{t}$ aggregates information about the downstream loss $\nabla J({\bm{\theta}}_{t})$ with respect to current training step and the LM’s training dynamics $\nabla^{2}L({\bm{\theta}}^{*}_{t},{\bm{\gamma}}^{*})$, from $T$ to $t$. As a result, ${\bm{\lambda}}^{*}_{t}$ summarizes the dynamics of LM pre-training (i.e., from future to the current). Since ${\bm{\gamma}}\in U$, Eq. (6) essentially suggests that $x_{n}$ with a higher $\sum_{t}{{\bm{\lambda}}^{*}_{t+1}}^{\top}\nabla l(x_{n},{\bm{\theta}}^{*}_{t})$ value should obtain a larger score in ${\bm{\gamma}}^{*}$, as shown in Figure 2 (Right). This indicates that the instances whose gradients align closely with the target vector ${\bm{\lambda}}^{*}_{t}$ should be selected. Note that the PMP conditions for data selection form a complete equation system, where ${\bm{\theta}}^{*}_{t}$, ${\bm{\lambda}}^{*}_{t}$, and ${\bm{\gamma}}^{*}$ are the solutions. In principle, by solving Eqs. (4)-(6) simultaneously, we can derive the optimal data quality scores, forming the foundation of our data selection framework PDS.
@@ 2.3.1 Data Quality Scores from PMP
Algorithm 1 solves the PMP conditions for data selection iteratively and returns the data quality scores ${\bm{\gamma}}^{*}$. We elaborate on the details of the algorithm and provide an efficient implementation.
Update of ${\bm{\gamma}}^{*}$: ${\bm{\gamma}}^{*}$ is supposed to be updated according to Eq. (6) with ${\bm{\theta}}^{*}_{t}$ and ${\bm{\lambda}}^{*}_{t}$ computed in the inner loops. Eq. (6) indicates that the new ${\bm{\gamma}}^{*}$ should be set as a one-hot vector, where the element corresponding to the highest $\sum_{t=0}^{T-1}{{\bm{\lambda}}^{*}_{t+1}}^{\top}\nabla l(x_{n},{\bm{\theta}}^{*}_{t})$ value is set to 1 and the others are set to 0. However, this “hard” update is unstable, as it causes training the LM with only one example in the upcoming outer loop iteration. Therefore, Algorithm 1 adopts a “soft” update, which increases ${\bm{\gamma}}^{*}$ by a value proportional to $\sum_{t=0}^{T-1}{\bm{\lambda}}_{t+1}^{\top}\nabla l(x_{n},{\bm{\theta}}_{t})$ and projects the updated ${\bm{\gamma}}^{*}$ back into $U$.
Figure 3: The PDS framework. We compute data quality scores ${\bm{\gamma}}^{*}$ on a proxy dataset $\mathcal{D}^{\mathrm{prx}}$ using Algorithm 1, which is derived from the Pontryagin’s Maximum Principle [63] (Section 2.3.1). After that, the data scorer learns to predict quality scores from instances, which then infers scores for a large corpus $\mathcal{D}$ (Section 2.3.2). Finally, a high-quality corpus $\mathcal{D}^{\prime}$ is selected based on the inferred scores to pre-train an LM (Section 2.3.3).
@@
         ${\bm{\theta}}_{t+1}\leftarrow{\bm{\theta}}_{t}-\eta\nabla L({\bm{\theta}}_{t},{\bm{\gamma}})$triangleright Corresponding to Eq. (4)
@@
         ${\bm{\lambda}}_{t}\leftarrow{\bm{\lambda}}_{t+1}+\nabla J({\bm{\theta}}_{t})-\eta\nabla^{2}L({\bm{\theta}}_{t},{\bm{\gamma}}){\bm{\lambda}}_{t+1}$triangleright Corresponding to Eq. (5)
@@
         $\gamma_{n}\leftarrow\gamma_{n}+\alpha\sum_{t=0}^{T-1}{\bm{\lambda}}_{t+1}^{\top}\nabla l(x_{n},{\bm{\theta}}_{t})$triangleright Corresponding to Eq. (6)
@@
where $\operatorname{PMP-Solver}$ refers to Algorithm 1. We also incorporate several practical optimization techniques to further reduce the computational overhead, as described in Appendix E.1.

### v2

@@ Abstract
This work investigates the selection of high-quality pre-training data from massive corpora to enhance LMs’ capabilities for downstream usage. We formulate data selection as a generalized Optimal Control problem, which can be solved theoretically by Pontryagin’s Maximum Principle (PMP), yielding a set of necessary conditions that characterize the relationship between optimal data selection and LM training dynamics. Based on these theoretical results, we introduce PMP-based Data Selection (PDS), a framework that approximates optimal data selection by solving the PMP conditions. In our experiments, we adopt PDS to select data from CommmonCrawl and show that the PDS-selected corpus accelerates the learning of LMs and constantly boosts their performance on a wide range of downstream tasks across various model sizes. Moreover, the benefits of PDS extend to ~400B models trained on ~10T tokens, as evidenced by the extrapolation of the test loss curves according to the Scaling Laws. PDS also improves data utilization when the pre-training data is limited, by reducing the data demand by 1.8 times, which helps mitigate the quick exhaustion of available web-crawled corpora. Our code, model, and data can be found at https://github.com/microsoft/LMOps/tree/main/data_selection.
@@
Figure 1: Scaling curves of average accuracy on 9 widely-used downstream tasks with respect to computation (a) and model sizes (b). We select pre-training corpora from the CommonCrawl and pre-train LMs on the selected data. PDS is compared with the Redpajama data cleaning pipeline (Together, 2023). The computation curves are calculated based on the training of a 1.7B LM.
@@ 1 Introduction
With the thriving of language models (LMs; Han et al., 2021; Bommasani et al., 2021), the role of data selection for pre-training becomes increasingly important, which aims at identifying valuable pre-training instances to accelerate model learning or improve downstream performance (Albalak et al., 2024). This focus enables researchers to explore the limit of LMs in the face of increasing training data demands (Brown et al., 2020; OpenAI, 2023; Team et al., 2023). It also helps to reduce the computational costs during pre-training (Sorscher et al., 2022), and addresses the potential limitations caused by available Internet data (Villalobos et al., 2022; Muennighoff et al., 2023). Without doubt, pre-training data selection is highly valuable for both research and industry sectors.
@@
Unlike previous works relying primarily on manually crafted heuristics (Tirumala et al., 2023; Xie et al., 2023), we connect data selection with classical Optimal Control theory (Lewis et al., 2012), where control variables in a dynamic system are optimized to achieve desired objectives. This mathematical formulation allows fine-grained white-box analysis of how the control variables drive a dynamic system from one state to another. In particular, by treating data selection as the control variables (i.e., whether a data point is included in pre-training), the LM pre-training process as the dynamic system, and the LM’s downstream performance as the objective, we leverage Pontryagin’s Maximum Principle (PMP; Pontryagin, 2018) to derive the necessary conditions for optimal data selection in theory. These results offer a rigorous, theory-driven alternative to the ad-hoc trial-and-error practices that currently dominate LM pre-training.
@@
Based on our theoretical results, we introduce PMP-based Data Selection (PDS), a framework that selects high-quality pre-training data at scale, by solving the equation system induced by the PMP conditions. Balancing effectiveness and efficiency, PDS first solves the equation system for the optimal data selection on a proxy dataset (e.g., 0.2B tokens), assigning a quality score to each instance based on its impact on downstream tasks. After that, a data scorer (e.g., 125M parameters) is trained to predict the quality scores and then infers scores on the target corpus (e.g., 50B tokens). Finally, the predicted scores guide data selection for pre-training LMs with various sizes (e.g., 1.7B parameters).
@@
Unlike previous pre-training data selection methods based on deduplication (Tirumala et al., 2023; Abbas et al., 2023), pattern matching (Xie et al., 2023), or single checkpoint performance (Engstrom et al., 2024), which are agnostic to the pre-training process of LMs, PDS exploits the highly dynamic nature of LM pre-training through the theoretical optimal control perspective. On the other hand, compared to methods that incorporate signals from the LM training process online (Yu et al., 2024; Wang et al., 2024), PDS operates offline, before the training begins, which avoids additional training-time computation overhead and allows for training LMs with arbitrary configurations while performing PDS only once. Furthermore, PDS only filters the training corpus, leaving highly optimized pre-training pipelines largely intact. Most importantly, PDS enjoys a strong theoretical basis, opening up the black box of understanding of individual data point impact on LM pre-training.
@@
In our experiments, we select data from the CommonCrawl with PDS using LIMA (Zhou et al., 2024) as the guide for downstream performance and pre-train LMs with 160M, 470M, 1B, and 1.7B parameters from scratch. Then, we test the LM’s zero-shot performance on tasks other than LIMA to examine the generalization of PDS. We observe around 2 times speed-up in pre-training on the 1.7B LM and constant improvement in downstream tasks and language modeling performance across all model sizes compared to state-of-the-art baselines. Extrapolating these results using the Scaling Law (Kaplan et al., 2020; Hoffmann et al., 2022), we show that the benefits remain consistent for ~400B LMs trained on ~15T tokens. Besides, PDS enhances data utilization in a data-constrained setting, reducing the pre-training data demand by 1.8 times, which is a critical advantage as the LM community is running out of data (Villalobos et al., 2022). We also conduct extensive analysis and ablation studies on the key factors of PDS to facilitate further research on data selection.
@@ 2.1 Problem Formulation
We study an LM parameterized with ${\bm{\theta}}\in{\mathbb{R}}^{N}$, pre-trained from scratch on a dataset $\mathcal{D}=\{x_{n}\}_{n=1}^{|\mathcal{D}|}$, over $T$ training steps. Data selection (Albalak et al., 2024) aims at identifying a subset $\mathcal{D}^{\prime}$ from $\mathcal{D}$, such that LMs trained on $\mathcal{D}^{\prime}$ achieve better performance, measured by a lower downstream loss $J({\bm{\theta}})$.
The pre-training process renders $J({\bm{\theta}})$ as a function of $\mathcal{D}^{\prime}$, which can be fully characterized by a data quality score vector ${\bm{\gamma}}=\left[\gamma_{1},\gamma_{2},\cdots,\gamma_{|\mathcal{D}|}\right]^{\top}$ in a $|\mathcal{D}|$-dimensional simplex $U$, where $U=\left\{\left.[\gamma_{1},\gamma_{2},\cdots,\gamma_{|\mathcal{D}|}]^{\top}\right|\sum_{n=1}^{|\mathcal{D}|}\gamma_{n}=1\text{ and }\gamma_{n}\geq 0\text{ for }1\leq n\leq{|\mathcal{D}|}\right\}$  [^1] Only the relative data quality is meaningful for data selection. Therefore, we ensure the sum of the quality scores to 1 to avoid the impact of their individual scales.. A higher quality score in ${\bm{\gamma}}$ indicates the corresponding instance is more helpful to reduce $J({\bm{\theta}})$, and thus the LM should learn more from the instance. Here, we focus on the independent importance of each example and provide a discussion on the dependence between each data point in Appendix E. This results in the following general pre-training loss, defined as the weighted sum of the per-instance loss by ${\bm{\gamma}}$:
@@
where $l(x_{n},{\bm{\theta}})=-\log p_{{\bm{\theta}}}(x_{n})$. The goal of data selection is thus to find ${\bm{\gamma}}$ that reduces the downstream loss $J({\bm{\theta}})$, and then select instances with the highest scores in ${\bm{\gamma}}$. For simplicity, we assume that the LM is trained using Gradient Decent (GD) for $0\leq t<T$, with the derivation under the Adam optimizer (Kingma & Ba, 2015) provided in Appendix C:
@@
Motivated by the literature on learned optimizers (Metz et al., 2020), we optimize ${\bm{\gamma}}$ by minimizing the area under the curve (AUC; Cortes & Mohri, 2003) of $J({\bm{\theta}}_{t})$, which is approximated by the cumulative sum of $J({\bm{\theta}}_{t})$ over the pre-training process:
@@
Intuitively, a lower AUC corresponds to faster convergence of the loss and improved final downstream performance. Unlike evaluating $J({\bm{\theta}}_{t})$ at individual time steps, the AUC captures the overall LM training dynamics. As shown in Appendix A, minimizing the AUC essentially enhances the constants in the LM’s Scaling Laws (Kaplan et al., 2020), leading to substantial improvements in LM learning.
@@ 2.2 Data Selection as Optimal Control
We recognize the optimization problem in Eq. (3) is analogous to a discrete optimal control problem (Lewis et al., 2012), where $J(\cdot)$ is the cost function, the model parameters ${\bm{\theta}}$ are the state variables evolving according to Eq. (2), and the data quality scores ${\bm{\gamma}}$ are the control variables to be optimized within $U$. This perspective makes it convenient for theoretically solving the data selection problem.
@@
Optimal control problems can be solved by a powerful tool known as Pontryagin’s Maximum Principle (PMP; Pontryagin, 2018), which provides a set of necessary conditions for the optimal control variables and their corresponding state variables (See Appendix B for its formal expression). However, standard PMP conditions allow the optimal control to vary over time, whereas in Eq. (3), the control variables ${\bm{\gamma}}$ are constrained to be time-invariant due to the offline nature of data selection in our setting. This typically makes the optimization problem more challenging due to the shrinking of feasible region. In the following, we present the PMP conditions for data selection under this constraint:
@@
We prove Theorem 2.1 in Appendix B using the standard PMP conditions and the Lagrange multiplier method, and provide an illustration for this theorem in Figure 2. Since standard PMP conditions and the Lagrange multiplier method are necessary conditions for the optimality, Theorem 2.1 is also a necessary condition for optimal data selection. By inspecting the PMP conditions for data selection, we can see that Eq. (4) ensures the LM parameters, trained with the optimal data quality scores, continue to evolve via GD. As illustrated in Figure 2 (Left), Eq. (5) defines ${\bm{\lambda}}^{*}_{t}$, a “target vector” suggesting the ideal gradient direction formed only by high-quality data points. In particular, ${\bm{\lambda}}^{*}_{t}$ aggregates information about the downstream loss $\nabla J({\bm{\theta}}_{t})$ with respect to current training step and the LM’s training dynamics $\nabla^{2}L({\bm{\theta}}^{*}_{t},{\bm{\gamma}}^{*})$, from $T$ to $t$. As a result, ${\bm{\lambda}}^{*}_{t}$ summarizes the dynamics of LM pre-training (i.e., from future to the current). Since ${\bm{\gamma}}\in U$, Eq. (6) essentially suggests that $x_{n}$ with a higher $\sum_{t}{{\bm{\lambda}}^{*}_{t+1}}^{\top}\nabla l(x_{n},{\bm{\theta}}^{*}_{t})$ value should obtain a larger score in ${\bm{\gamma}}^{*}$, as shown in Figure 2 (Right). This indicates that the instances whose gradients align closely with the target vector ${\bm{\lambda}}^{*}_{t}$ should be selected. Note that the PMP conditions for data selection form a complete equation system, where ${\bm{\theta}}^{*}_{t}$, ${\bm{\lambda}}^{*}_{t}$, and ${\bm{\gamma}}^{*}$ are the solutions. In principle, by solving Eqs. (4)-(6) simultaneously, we can derive the optimal data quality scores, forming the foundation of our data selection framework PDS.
@@ 2.3.1 Data Quality Scores from PMP
         ${\bm{\theta}}_{t+1}\leftarrow{\bm{\theta}}_{t}-\eta\nabla L({\bm{\theta}}_{t},{\bm{\gamma}})$triangleright Eq. (4)
@@
         ${\bm{\lambda}}_{t}\leftarrow{\bm{\lambda}}_{t+1}+\nabla J({\bm{\theta}}_{t})-\eta\nabla^{2}L({\bm{\theta}}_{t},{\bm{\gamma}}){\bm{\lambda}}_{t+1}$triangleright Eq. (5)
@@
         $\gamma_{n}\leftarrow\gamma_{n}+\alpha\sum_{t=0}^{T-1}{\bm{\lambda}}_{t+1}^{\top}\nabla l(x_{n},{\bm{\theta}}_{t})$triangleright Eq. (6)
@@
Algorithm 1 solves the PMP conditions for data selection iteratively and returns the data quality scores ${\bm{\gamma}}^{*}$.
Update of ${\bm{\gamma}}^{*}$: ${\bm{\gamma}}^{*}$ is supposed to be updated according to Eq. (6) with ${\bm{\theta}}^{*}_{t}$ and ${\bm{\lambda}}^{*}_{t}$ computed in the inner loops. Eq. (6) indicates that the new ${\bm{\gamma}}^{*}$ should be set as a one-hot vector, where the element corresponding to the highest $\sum_{t=0}^{T-1}{{\bm{\lambda}}^{*}_{t+1}}^{\top}\nabla l(x_{n},{\bm{\theta}}^{*}_{t})$ value is set to 1 and the others are set to 0. However, this “hard” update is unstable, as it causes training the LM with only one example in the upcoming outer loop iteration  [^2] This does not imply that one single example is optimal for data selection due to the necessity but not sufficiency of Theorem 2.1. See Appendix F for a detailed discussion. Therefore, Algorithm 1 adopts a “soft” update, which increases ${\bm{\gamma}}^{*}$ by a value proportional to $\sum_{t=0}^{T-1}{\bm{\lambda}}_{t+1}^{\top}\nabla l(x_{n},{\bm{\theta}}_{t})$ and projects the updated ${\bm{\gamma}}^{*}$ back into $U$.
Figure 3: The PDS framework. We compute data quality scores ${\bm{\gamma}}^{*}$ on a proxy dataset $\mathcal{D}^{\mathrm{prx}}$ using Algorithm 1, which is derived from the Pontryagin’s Maximum Principle (Pontryagin, 2018) (Section 2.3.1). After that, the data scorer learns to predict quality scores from instances, which then infers scores for a large corpus $\mathcal{D}$ (Section 2.3.2). Finally, a high-quality corpus $\mathcal{D}^{\prime}$ is selected based on the inferred scores to pre-train an LM (Section 2.3.3).
@@
where $\operatorname{PMP-Solver}$ refers to Algorithm 1. We also incorporate several practical optimization techniques to further reduce the computational overhead, as described in Appendix G.1.